1. A 1.60 m long steel piano wire has a diameter of 0.20 cm. How great is the tension in the wire if it stretches 0.25 cm when tightened?
1
Expert's answer
2020-07-14T08:51:53-0400

The relation between the tension and stretch is given by the Hooke's law:


"\sigma = Y\varepsilon"

where "\sigma = \dfrac{F}{A}" is the stress (tension per unit area), "\varepsilon = \dfrac{\Delta L}{L}" is the relative stretch ("\Delta L = 0.25\times10^{-2}m" and "L = 1.6m" ) and "Y = 200\times10^9Pa" is the Young's modulus for steel.The cross-sectional area of the string is "A = \pi d^2/4"

Combining it all together, obtain:

"\dfrac{F}{A} = Y\dfrac{\Delta L}{L}"

Expressing the tension, obtain:


"F = \dfrac{AY\Delta L}{L} = \dfrac{\pi d^2Y\Delta L}{4L}= \dfrac{\pi 0.002^2\cdot 200\cdot 10^9 \cdot 0.0025}{1.6} \approx 3.93\times 10^3 N"

Answer. 3.93*10^3 N.